Calculate the activity coefficient, ?, of Cu2 when the ionic strength of the solution, ? , is 0.078 M:

(a) by linear interpolation of the data in the table to the right;

yCu2+ = 0.440

(b) by using the extended Debye-Huckel equation at 25°C, where the ion size is 600 pm.

yCu2+ =

answer to part a is correct.

ionic strength (\mu,M) activity coeffient \gamma Cu2+
0.001 0.870
0.005 0.749
0.01 0.675
0.05 0.485
0.1 0.405
Answer #1

For b.

The extended Debye-Hückel equation

Log y = (-0.51 · z2 · u1/2) / (1+ 3.3 · a · u1/2 )

          = (-0.51 · 22 · 0.0781/2) / (1+ 3.3 · 0.6 · 0.0781/2 )

          = - 0.01241 / 1.01204

          = - 0.01226

        y = 10- 0.01226 = 0.972

Note:

For the effective diameter a, use the value in nm for calculation:

         600 pm = 0.600 nm